OpenStudy (migitmack):

Solve v=(1)/(3)BH for b

OpenStudy (anonymous):

Divide both the sides H first and multiply both the sides by 3..

OpenStudy (anonymous):

\[\frac{V}{H} = \frac{1}{3}B \implies B = \frac{3V}{H}\]

OpenStudy (anonymous):

\[V= 1/3*B*H\] K so for this, we will multiply the fraction by its reciprocal (opposite) The opposite of 1/3 is 3, so we multiply 3 on both sides to cancel the fraction out. \[3*V= (1/3*3/1)*B*H\] This now gives us \[3V=B*H\] Now we divide both sides by 3H and get \[3V/3=(B*H)/H\] We finally get, \[3V/H=B\]